OpenStudy (anonymous):

I need help with this last question. It is still Subtracting Polynomials. Subtract 10x^3+4x^2+(-7) from -2x^3+9x^3+(-2x)

OpenStudy (savvy):

did you try using parenthesis....

OpenStudy (anonymous):

Yes, I also tried doing what you did, but it keeps on saying wrong answer. Dx I don't know why, but it tells me it needs to be a positive answer.

OpenStudy (savvy):

ok type here wat u did...

OpenStudy (anonymous):

(10x^3+4x^2-7)-(-2x^3+9x^2-2x) (10x^3+4x^2-7)+(2x^3-9x^2+2x) 10x^3+4x^2-7+2x^3-9x^2+2x 12x^3-5x^2-7+2x

OpenStudy (savvy):

dude the question says "SUBTRACT 10x^3+4x^2+(-7) FROM -2x^3+9x^3+(-2x)"....you've done the opposite....

OpenStudy (anonymous):

D: Oh...

OpenStudy (anonymous):

So then it would be, (10x^3+4x^2-7)-(-2x^3+9x^2-2x) 10x^3+4x^2-7+2x^3-9x^2+2x 12x^3-5x^2-7+2x

OpenStudy (savvy):

what is the difference...??? i meant that you have to do this...: (-2x^3+9x^2-2x) - (10x^3+4x^2-7)

OpenStudy (anonymous):

And then I combine like terms?

OpenStudy (savvy):

yeah....

OpenStudy (anonymous):

Thank you!

OpenStudy (campbell_st):

the easy way to do this question is to simply multiply \[10x^3 + 4x^2 -7 \] by negative 1 and add \[(10x^3 + 4x^2 - 7) \times -1 = -10x^3 - 4x^2 + 7\] now you have \[-2x^3 +9x^2 -2x -10x^3 - 4x^2 + 7 = -12x^3 +5x^2 -2x +7\]